36-303 TEAM WORKING AGREEMENT (TWA) ASSIGNMENT Due Tuesday Feb 23, 2010, in email and \underline{ON PAPER}

Your team will create a Team Working Agreement (TWA) in item 8 below. This will be a written contract that your team creates to describe what individual responsibilities will be and how your team will handle problems. Problems should be handled primarily in your team. As a last resort, you may choose to go to someone else, for example the instructor or the TA for help. Your TWA should specify when you would do that.

You will be turning in, for credit, your Team Goals Statement (item 7 below) and Team Working Agreement (item 8 below), signed by all team members. In order for this to be a success, it is crucial that you spend the time needed to work through the first 6 exercises on your own—and as a team—in preparation for your goals statement and working agreement.

Complete the following activities individually:

1. Think about teams you have been on in all areas of your life: sports, clubs, music, drama, church, family, school courses, etc. Name one of these teams that you think was successful, and one team that you think was unsuccessful. What differences do you see between the successful and unsuccessful teams? Report on you how felt on each of the teams as well as the reasons you think the teams were successful or unsuccessful.

2. Throughout the semester in this course you will be working in a project team. Imagine that this team is an ideal team. What characteristics will it possess?

3. What thoughts do you have about working on teams?

Share your responses to Questions 1–3 with your teammates in order to begin to get to know each other better. Compile your results. Then work through the dilemmas below with your team.

4. Dilemma One:

You are working in a team of three or four people. Responsibilities have been assigned and agreed upon by all. At your next meeting everyone is prepared but Stephanie.

(a) What will you say to Stephanie?

(b) Stephanie explains to the team that she was not feeling well the night before and was unable to complete her responsibility. What is your response?

(c) What does your team do to continue progress on your project?

(d) How does your team document this discussion and the actions taken?

The next day Stephanie comes prepared and the team progresses well. A week later on the day your team was going to put together your project, Michael forgets to bring the materials he agreed to bring.

(e) How does your team respond?

(f) Michael had a calculus exam that he had to pass to avoid failing the class; does this affect your actions?

(g) How will you make the progress you need to make?

The next day is Friday, Michael brings his materials, but Stephanie is absent. Your project is due on Monday.

(h) How does your team respond?

(i) Stephanie had to leave Thursday to go home for the weekend. Her home is 7 hours driving and her parents said she had to leave Thursday. Does this affect your decision?

(j) What can you do as a team to plan for delays?

5. Dilemma Two:

Suppose three of the four of you regularly attend meeting but one member is always absent, making excuses, or appears to have other priorities, or does not respond to repeated emails, calls etc.

(a) What do you plan to say if and when you confront the fourth member?

(b) At what point do you go to someone else for help with your dilemma?

(c) How will you document these events so that at the end of the semester you are not simply accusing each other?

6. Dilemma Three:

You are working in a team with John. He claims to always know what to do and he is always sure his design will work. Because he is so sure, he doesn’t listen to other people’s suggestions, or he puts others’ suggestions down. For example, if someone tries to edit a document for the team, John always edits it back to the way he likes it to read.

(a) How will you work with John in the team?

(b) Under what circumstances will you ask someone else for help?

(c) What is your plan to avoid these two pitfalls:

• Everyone does what John says, ideas are unchallenged and the confidence proves unfounded.

• The team breaks into warring camps with some taking John’s side, other standing in opposition.

7. Team Goals Statement Create a statement of your team’s goals in this project. The statement should be three to four sentences and should neither be simplistic nor facetious. For example, “Get a good grade” is trite. Everyone wants to get a “good grade” but for some that is an “A”, others it is anything above a “D”. Everyone wants to achieve the 100-point performance level, but not everyone cares about the competition. Most partners don’t mind “working on it” per se, but some are willing to put in whatever time it takes to achieve something they are proud of, others have a strict time budget they are not willing to exceed.

8. Team Working Agreement

In light of your responses to the previous questions, write a TWA that enhances the chances that your team will achieve its mission. Be sure your TWA addresses the concerns each of you has about working on a team, incorporates your strategies for dealing with the dilemmas, and specifies the circumstances under which each of you will leave the team and work on your own or with a subset of the original team.

Answer the following questions (a)–(l) in your TWA:

(a) What times in the week can you get together as a team, and where?

(b) How will you contact each other about meetings?

(c) How much advance notice will you give before cancelling meetings?

(d) Will you allow future amendments to your TWA? If so, how will this be done?

(e) What do you expect each member to contribute to your team?

(f) What behaviors are acceptable in team meetings?

(g) How will team members deal with disagreements?

(h) How will the team deal with absences?

(i) Under what conditions will your team choose to go to someone else for help?

(j) What other policies will you establish to address situations similar to the dilemmas above?

Suppose that, in spite of all your planning, worse comes to worst and it appears that one or more members of the team are not doing their jobs, or their actions are preventing others from doing their jobs.

(k) Under what conditions will you give up trying to get a team member to contribute and instead seek to document this non-performance for submission to the instructor? How many missed or partially missed meetings? How long a delay in returning email? How many unilateral decisions? Etc.

(l) What warnings will you give and what mechanism do all agree is suitable for delivery of a warning?

Type up your Team Goals Statement and your Team Working Agreement. Attach all the names of the team members to this final document, and have each team member sign and date the document above his/her name. Turn in:

• Electronic copy of this final document (missing signatures if need be), as a pdf or msword attachment in email to me.

• Signed paper copy of the final document. Make sure that each team member has a copy of the final document as well. This is your contract with the team, for the rest of the semester.

Both email and hardcopy of this final document is due Tuesday, Feb 23, 2008, by 5:00pm. This will count as one quiz grade (100 pts) in the class.